A 2 kg rectangular thin plate has a length of 0.5 m and a width of 0.20. the moment of inertia


of the plate is 0.29 kg-m2


. Where is the location of the axis of rotation?



1
Expert's answer
2022-04-19T16:32:00-0400

"I_0=\frac{m(a^2+b^2)}{12}=\frac I6,"

"I=6I_0=I_0+md^2,"

"md^2=5I_0=\frac 56I,"

"d=\sqrt\frac{5I}{6m}=0.35~m."